OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)? OPTIONS: A) y – 5 = 3(x – 1) B) y – 2 = –3(x – 0) C) y + 2 = 3(x + 1) D) y + 5 = –3(x + 1)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(0\quad ,&2)\quad &(1\quad ,&5) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (jdoe0001):

find the slope first

OpenStudy (jdoe0001):

then use that in the "point-slope form" of \(\bf y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

2-5/1-0= -3/1=-3

OpenStudy (anonymous):

y-5=-3(x-0)

OpenStudy (jdoe0001):

well. \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(0\quad ,&2)\quad &(1\quad ,&5) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{5-2}{1-0}\)

OpenStudy (jdoe0001):

so that gives you a slope of " 3 "

OpenStudy (jdoe0001):

then pick either point and plug it in the \(\bf y-y_1=m(x-x_1)\)

OpenStudy (jdoe0001):

then solve for "y"

OpenStudy (anonymous):

y-5=3(x-0) y-5=3x+0 y=3x+5 @jdoe0001 ?????

OpenStudy (jdoe0001):

hmmm.... well, you don't have a point of ( 0, 5) for one to use y-5=3(x-0)

OpenStudy (jdoe0001):

(0, 2) and (1, 5) <--- either point will do for the "point-slope" form though

OpenStudy (anonymous):

so is A correct? @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

alright thanks @jdoe0001

OpenStudy (jdoe0001):

yw